\begin{thebibliography}{10}






\bibitem{cacdan}
F. Cacciafesta and P. D'Ancona.
\newblock Endpoint estimates and global existence for the nonlinear Dirac equation with potential.
\newblock  {\em J. Diff. Eq.}, 254(5):2233--2260, 2013. 





%\bibitem{CaffarelliKohnNirenberg84-a}
%Luis~A. Caffarelli, Robert Kohn, and Louis Nirenberg.
%\newblock First order interpolation inequalities with weights.
%\newblock {\em Compositio Math.}, 53(3):259--275, 1984.


\bibitem{coiffeff}
R. R. Coifman and C. Fefferman, C.
\newblock {Weighted norm inequalities for maximal functions and singular integrals.}
\newblock {\em Studia Math.}, 51:241--�250, 1974. 

\bibitem{Cordoba}
A. C\'{o}rdoba.
\newblock Singular integrals and maximal functions: the disk multiplier revisited.
\newblock {\em Adv. Math.}, 290: 208--235, 2016.


\bibitem{CordobaFefferman}
A. C\'{o}rdoba and C. Fefferman.
\newblock A weighted norm inequality for singular integrals.
\newblock {\em Studia Math.}, 57(1):97--101, 1976.






%\bibitem{CKN} 
%L. Caffarelli, R. Kohn and L. Nirenberg. 
%\newblock Partial regularity of suitable weak solutions of the Navier--Stokes equations. 
%\newblock {\em Comm. Pure Appl. Math.}, 35:771--831, 1982.
%
%
%
%\bibitem{ChoOzawa09-a}
%Y. Cho and T. Ozawa.
%\newblock Sobolev inequalities with symmetry.
%\newblock {\em Commun. Contemp. Math.}, 11(3):355--365, 2009.



\bibitem{DL} 
P. D'Ancona and R. Luc\`a.
\newblock Stein-Weiss and Caffarelli-Kohn-Nirenberg inequalities with higher
angular integrability.
\newblock {\em J. Math. Anal. App.}, 388(2):1061--1079, 2012.


\bibitem{DLNS} 
P. D'Ancona and R. Luc\`a.
\newblock On the regularity set and angular integrability for the Navier--Stokes equation.
\newblock ArXiv:1501.07780.

%
%
%%\bibitem{DenapoliDrelichmanDuran09-a}
%%P.~L. De~N{\'a}poli, I. Drelichman and R.~G. Dur{\'a}n.
%%\newblock Radial solutions for {H}amiltonian elliptic systems with weights.
%%\newblock {\em Adv. Nonlinear Stud.}, 9(3):579--593, 2009.
%
%
%\bibitem{DenapoliDrelichman09-aBis}
% P.~L. De~N{\'a}poli and I. Drelichman.
%\newblock Weighted convolution inequalities for radial functions. 
%\newblock {\em Ann. Mat. Pur. Appl.}, to appear. 
%
%
%
%
%\bibitem{DenapoliDrelichmanDuran10-a}
% P.~L. De~N{\'a}poli, I. Drelichman and R.~G. Dur{\'a}n.
%\newblock Improved Caffarelli-Kohn-Nirenberg and trace inequalities for radial functions. 
%\newblock {\em  Comm. Pure Appl. Anal.}, 11(5):1629--1642, 2012.  
%
%

\bibitem{DenapoliDrelichmanDuran11-a}
P.~L. De~N{\'a}poli, I. Drelichman and R.~G. Dur{\'a}n.
\newblock On weighted inequalities for fractional integrals of radial functions.
\newblock {\em Illinois J. Math.}, 55:575--587, 2011.



%\bibitem{Esc}
%L. Escauriaza, G. Seregin and V. Sverak
%\newblock Backward uniqueness for parabolic equations.
%\newblock {\em Arch. Ration. Mech. Anal.}, 169:147--157, 2003.




\bibitem{FangWang08-a}
D. Fang and C. Wang.
\newblock Weighted {S}trichartz estimates with angular regularity and their applications.
\newblock {\em Forum Math.}, 23:181--205, 2011.



\bibitem{Grafakos08-a}
Loukas Grafakos.
\newblock {\em Classical {F}ourier analysis}, volume 249 of {\em Graduate Texts
  in Mathematics}.
\newblock Springer, New York, second edition, 2008.


\bibitem{Ren}
R. Luc{\`a}.
\newblock Regularity criteria with angular integrability for the Navier--Stokes equation.
\newblock {\em Nonlinear Anal.}, 105:24--40, 2014. 


%\bibitem{Giga} 
%Y. Giga. 
%\newblock Solutions for semilinear parabolic equations in $L^{p}$ and regularity of weak solutions of the Navier--Stokes system. 
%\newblock {\em J. Diff. Eq.}, 62:186--212, 1986.

%
%
%
%\bibitem{Hopf} 
%E. Hopf. 
%\newblock Uber die Anfanqswertaufgabe f\"{u}r die hydrodynamischen Grundgleichungen. 
%\newblock {\em Math. Nachr.}, 4:213--231, 1951.
%
%
%
%
%\bibitem{Ren} 
%R. Luc\`a.
%\newblock ?????????.
%\newblock {\em ?????}, ?????.
%
%
%
%\bibitem{Kat}
%T. Kato. 
%\newblock Strong $L^{p}$-solutions of the Navier--Stokes equation in $\mathbb{R}^{n}$, with applications to weak solutions. 
%\newblock {\em Math. Z.}, 187: 471--480, 1984.
%
%
%
%
%\bibitem{Kuk1}
%I. Kukavica.
%\newblock On partial regularity for the Navier--Stokes equations.
%\newblock {\em Discrete Contin. Dyn. Syst.}, 21:717--728, 2008.
%
%
%\bibitem{Kuk2}
%I. Kukavica.
%\newblock On regularity for the Navier--Stokes equations in Morrey spaces. 
%\newblock {\ Discrete Contin. Dyn. Syst.} 26(4):1319--1328, 2010. 
%
%
%
%
%
%\bibitem{Lem}
%P. G. Lemari{\'e}-Rieusset.
%\newblock Recent developments in the Navier--Stokes problem.
%\newblock CHAPMAN AND HALL/CRC. Research Notes in Mathematics Series 431, 2002.
%


%
%\bibitem{Ler} 
%J. Leray. 
%\newblock Sur le mouvement d'un liquide visqueux emplissant l'espace. 
%\newblock {\em Acta Math.}, 63:193--248, 1934.





\bibitem{MachiharaNakamuraNakanishi05-a}
S. Machihara, M. Nakamura, K. Nakanishi, and T. Ozawa.
\newblock Endpoint {S}trichartz estimates and global solutions for the
  nonlinear {D}irac equation.
\newblock {\em J. Funct. Anal.}, 219(1):1--20, 2005.




\bibitem{Rogers}
T. Ozawa and K. M. Rogers
\newblock Sharp Morawetz estimates.
\newblock {\em J. Anal. Math.}, 121:163--175, 2013.


%
%
%
%\bibitem{Rubin83-a}
%B. Rubin.
%\newblock One-dimensional representation, inversion and certain 
%properties of Riesz potentials of radial functions. (russian).
%\newblock {\em Mat. Zametki}, 34(4):521---533, 1983.
%\newblock English translation: Math. Notes 34(3-4):751--757, 1983.



%\bibitem{S1}
%V. Scheffer.
%\newblock Partial regularity of solutions to the Navier-Stokes equations.
%\newblock {\em Pacific J. Math.}, 66(2):535---5552, 1976.
%\newblock English translation: {\em Math. Notes}, 34(3-4):751--757, 1983.
 


%\bibitem{S2}
%V. Scheffer.
%\newblock Hausdroff measure and the Navier-Stokes equations.
%\newblock {\em Comm. Math. Phys.}, 55(2):97---112, 1977.
%\newblock English translation: {\em Math. Notes}, 34(3-4):751--757, 1983.

%
%
%
%\bibitem{Ser} 
%J. Serrin. 
%\newblock On the interior regularity of weak solutions of the Navier--Stokes equations. 
%\newblock {\em Arch. Ration. Mech. Anal.}, 9:187--195, 1962.  
%
%\bibitem{Sohr} 
%H. Sohr. 
%\newblock Zur Regularit\"{a}tstheorie der instationaren Gleichungen von Navier--Stokes.
%\newblock {\em Math. Z.}, 184:339--375, 1983.  


\bibitem{Stein}
E. M. Stein.
\newblock Note on singular integrals.
\newblock {\em Proc. Am. Math. Soc.}, 8:250--254, 1957.


%
%
%\bibitem{Struwe} 
%M. Struwe.
%\newblock On partial regularity results for the Navier--Stokes equations.
%\newblock {\em Commun. Pure Appl. Math.} 41:437--458, 1988.
%
%\bibitem{Tay}
%M. E. Taylor 
%\newblock Analysis on Morrey spaces and applications to Navier--Stokes and other evolution equations. 
%\newblock {\ Comm. Part. Diff. Eq.}, 17(9-10):1407--1456, 1992. 
%
%
%
%\bibitem{YongZhou} 
%Y. Zhou. 
%\newblock Weighted regularity criteria for the three-dimensional Navier--Stokes equations. \newblock {\em Proc. Royal Soc. Edin.}, 139A:661--671, 2009.


%\bibitem{GinibreVelo95-b}
%Jean Ginibre and Giorgio Velo.
%\newblock Generalized {S}trichartz inequalities for the wave equation.
%\newblock {\em J. Funct. Anal.}, 133(1):50--68, 1995.


%\bibitem{HidanoKurokawa08-a}
%Kunio Hidano and Yuki Kurokawa.
%\newblock Weighted {HLS} inequalities for radial functions and {S}trichartz
%  estimates for wave and {S}chr{\"o}dinger equations.
%\newblock {\em Illinois J. Math.}, 52(2):365--388, 2008.

%\bibitem{JiangWangYu10-a}
%Chengbo~Wang Jin-Cheng~Jiang and Xin Yu.
%\newblock Generalized and weighted strichartz estimates.
%\newblock 2010.

%\bibitem{KeelTao98-a}
%Markus Keel and Terence Tao.
%\newblock Endpoint {S}trichartz estimates.
%\newblock {\em Amer. J. Math.}, 120(5):955--980, 1998.


%\bibitem{Ni82-a}
%Wei~Ming Ni.
%\newblock A nonlinear {D}irichlet problem on the unit ball and its
%  applications.
%\newblock {\em Indiana Univ. Math. J.}, 31(6):801--807, 1982.

%\bibitem{SawyerWheeden92-a}
%Eric~T. Sawyer and Richard~L. Wheeden.
%\newblock Weighted inequalities for fractional integrals on {E}uclidean and
%  homogeneous spaces.
%\newblock {\em Amer. J. Math.}, 114(4):813--874, 1992.

%\bibitem{SickelSkrzypczak00-a}
%Winfried Sickel and Leszek Skrzypczak.
%\newblock Radial subspaces of {B}esov and {L}izorkin-{T}riebel classes:
%  extended {S}trauss lemma and compactness of embeddings.
%\newblock {\em J. Fourier Anal. Appl.}, 6(6):639--662, 2000.


%\bibitem{SteinWeiss58-b}
%Elias~M. Stein and Guido Weiss.
%\newblock Fractional integrals on {$n$}-dimensional {E}uclidean space.
%\newblock {\em J. Math. Mech.}, 7:503--514, 1958.

\bibitem{Sterbenz05-a}
J. Sterbenz.
\newblock Angular regularity and {S}trichartz estimates for the wave equation.
\newblock {\em Int. Math. Res. Not.}, 4:187--231, 2005.
\newblock With an appendix by Igor Rodnianski.



\bibitem{TaoSphere}
T. Tao.  
\newblock Spherically averaged endpoint Strichartz estimates for the two-dimensional Schr\"dinger equation. 
\newblock {\em Comm. Partial Differential Equations} 25(7-8):1471--1485, 2000.

%
%\bibitem{WW}
%W. von Wahl. 
%\newblock Regularity of weak solutions of the Navier--Stokes equations. In Proc. 1983
%Summer Inst. on Nonlinear Functional Analysis and Applications, Proceedings of Symposia
%in Pure Mathematics, vol. 45, pp. 497--503 (Providence, RI: American Mathematical Society,
%1989).

%\bibitem{Strauss77-a}
%Walter~A. Strauss.
%\newblock Existence of solitary waves in higher dimensions.
%\newblock {\em Comm. Math. Phys.}, 55(2):149--162, 1977.

%\bibitem{Vilela01-a}
%Maria~Cruz Vilela.
%\newblock Regularity of solutions to the free {S}chr{\"o}dinger equation with
%  radial initial data.
%\newblock {\em Illinois J. Math.}, 45(2):361--370, 2001.

\end{thebibliography}

\documentclass[10pt,oneside,a4paper,reqno]{amsart}  

\usepackage{amssymb,graphicx,xcolor}
\usepackage[colorlinks=true, linkcolor=magenta, citecolor=cyan, urlcolor=blue]{hyperref}
\usepackage{bbm}

\usepackage{graphicx}
\usepackage{mathrsfs}

\numberwithin{equation}{section}\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{remark}
\newtheorem{remark}{Remark}[section]
\theoremstyle{definition}

\newcommand{\p}{\widetilde{p}}

\newcommand{\Rn}{\mathbb{R}^{n}}
\newcommand{\Rpiu}{\mathbb{R}^{+}}

\title{Singular integrals with angular integrability}
\date{}    

\author{Federico Cacciafesta}
\address{Federico Cacciafesta: 
SAPIENZA --- Universit\`a di Roma,
Dipartimento di Matematica, 
Piazzale A.~Moro 2, I-00185 Roma, Italy}
\email{cacciafe@mat.uniroma1.it}

\author{Renato Luc\`a}
\address{Renato Luc\`a: Instituto de Ciencias Matem\'aticas CSIC-UAM-UC3M-UCM, Madrid, 28049, Spain.}
\email{renato.luca@icmat.es}

\subjclass[2010]{42B37, 42B20}
\keywords{Singular integrals. Angular integrability}

\begin{document}

 \begin{abstract}
In this note we prove a class of sharp inequalities for singular integral operators in weighted Lebesgue spaces with angular integrability.

\end{abstract}

\maketitle

\section{Introduction}

We consider singular integral operators
\begin{equation}\label{singint}
T f (x):=
\mathrm{P.V.}
\int_{\mathbb{R}^{n}} f(x-y) K(y) \ dy,
\end{equation}
where the kernel $K$ satisfies the following conditions
$$
|y|^{n} |K| \leq C,
\quad
|y|^{n+1} | \nabla K | \leq C,
\quad
| \widehat{K} | \leq C.
$$
Here $C >0 $ is a constant and $\, \widehat{} \,$ denotes the Fourier transform.
The main example we have in mind is the directional Riesz transform, which corresponds to the choice 
$K(y) := |y|^{-(n+1)} y \cdot \theta$, $\theta \in \mathbb{S}^{n-1}$.

The study of the boundedness of these operators in weighted Lebesgue spaces $L^p(w(x)dx)$,  for $1 < p <\infty$ and $0 < w \in L^{1}_{loc}(\Rn)$, is a classical problem in harmonic analysis: in particular, Stein \cite{Stein} proved it for the (sharp range of) homogeneous weights $w(x)=|x|^{\alpha p}$, $-n/p < \alpha < n- n/p$. The result was later extended by Coifman and Fefferman \cite{coiffeff} to any $A_p$ weight.
%

While the weighted $L^{p}$-theory has been extensively studied, less is known in the case of Lebesgue norms with different integrability in the radial and angular directions, namely

\begin{equation}\label{MixedNorms}
\begin{array}{lcl}
\displaystyle
  \|f\|_{L^{p}_{|x|}L^{\p}_{\theta}}& := &
\displaystyle
 \left(
    \int_{0}^{+\infty}
    \|f(\rho \, \cdot \, )\|^{p}_{L^{\p}(\mathbb{S}^{n-1})}
    \rho^{n-1}d \rho
 \right)^{\frac1p}. 
\end{array}
\end{equation}

These mixed radial-angular spaces have been successfully used in recent years to improve several results in the framework of 
partial differential equations; see e.g. \cite{cacdan, FangWang08-a, Ren, MachiharaNakamuraNakanishi05-a, Rogers, Sterbenz05-a, TaoSphere}. 
Notice that when $p=\p$ the norms reduce to the usual $L^{p}$ norms. 
Notice also that, neglecting the constants, they are increasing in $\p$, and that they behave as the $L^{p}$ norms under homogeneous rescaling, namely
$f(\cdot) \to f(\lambda \, \cdot \, )$, $\lambda > 0$.

In a recent paper A. C\'{o}rdoba \cite{Cordoba} proved, among the other things, 
the 
$L^{p}_{|x|}L^{2}_{\theta}$ boundedness for operators of the form (\ref{singint}).
Here we give an extension of this result to the weighted setting.

\begin{theorem}\label{OurSteinWeiss}
Let $n \geq 2$, $1 < p < \infty$, $1 < \p < \infty$ and $-n/p < \alpha < n- n/p$. 
Then
\begin{equation}\label{eq:OurSteinWeiss}
\| |x|^{\alpha} T \phi \|_{L^{p}_{|x|}L^{\p}_{\theta}} 
\leq C \| |x|^{\alpha} \phi \|_{L^{p}_{|x|}L^{\p}_{\theta}},
\end{equation}  
where $C$ is a constant depending only on $\alpha, p, \p, n$.
\end{theorem}

Let us point out that the case $\alpha =0$, $1 < \p < \infty$ in inequality (\ref{eq:OurSteinWeiss}) may be deduced by the application of C\'{o}rdoba's argument; see \cite[Theorem 2.6]{DLNS}. 
Therefore the novelty of Theorem \ref{OurSteinWeiss} is in that it covers all the possible homogeneous weights of the kind $|x|^{\alpha p}$.

\begin{remark}
Condition $-n/p < \alpha$ turns to be necessary by testing the inequality on functions $\phi$ such that $\phi=0$ and $T\phi>0$ in a neighborhood of the origin. On the other hand, condition $\alpha < n - n/p$ turns to be necessary for the same reason by considering the dual inequality.
\end{remark}

\begin{remark}
One of the estimates (\ref{eq:OurSteinWeiss}) has been used in \cite[Theorem 1.5]{DLNS} to deduce informations
about the regularity of weak solutions of the $3d$ Navier--Stokes problem with initial velocities
satisfying good angular integrability properties.
\end{remark}

We write $A \lesssim B$ if $A \leq C B$ with a constant $C$ depending only on $\alpha, p, \p, n$. We write $A \simeq B$ if both $A \lesssim B$ and $B \lesssim A$.

\section{Proof}
\setcounter{equation}{0}

We know by \cite[Theorem 2.6]{DLNS} that inequality (\ref{eq:OurSteinWeiss}) is true in the case $\alpha =0$, that is
\begin{equation}\label{Eq:NONweightedSW}
\| T g \|_{L^{p}_{|x|}L^{\p}_{\theta}} \lesssim \| g \|_{L^{p}_{|x|}L^{\p}_{\theta}}.
\end{equation}

Following Stein \cite{Stein}, we now show that the weighted case (\ref{eq:OurSteinWeiss}) can be then deduced by the unweighted one. The next lemma represents the core of the proof.

\begin{lemma}\label{SteinMainLemma}
Let $n \geq 2$, $1 < p < \infty$, $1\leq \p \leq \infty$, $-n/p < \alpha < n - n/p$ and
\begin{equation}\label{eq:SteinWeissKernel}
F(x,y) := \frac{|1- (|x|/|y|)^{\alpha}|}{|x-y|^{n}},
\end{equation}
then
\begin{equation}\label{Eq:SWInequalityGeneralized}
\bigg\| \int_{\Rn} F(x,y) \phi(y) dy  \bigg\|_{L^{p}_{|x|}L^{\p}_{\theta}}
\leq C 
\| \phi \|_{L^{p}_{|x|}L^{\p}_{\theta}}.
\end{equation}
\end{lemma}

Assume indeed this has been proved and
first apply inequality \eqref{Eq:NONweightedSW} with the choice $g:= |\cdot|^{\alpha} f$ to have
\begin{equation}\label{SWQuasiI}
\| T (|x|^{\alpha}f) \|_{L^{p}_{|x|}L^{\p}_{\theta}} \lesssim \| |x|^{\alpha} f \|_{L^{p}_{|x|}L^{\p}_{\theta}}.
\end{equation}
Then notice that  
\begin{equation}\label{SWQuasiII}
\begin{split}
& \big| T(|x|^{\alpha} f)  -   |x|^{\alpha} T f \big| 
 \leq
 \int_{\Rn}  \big| K(x-y) ( |y|^{\alpha} - |x|^{\alpha}) f(y) \big|  dy  
\\ 
& \lesssim
 \int_{\Rn}    \frac{| |y|^{\alpha} - |x|^{\alpha} |}{|x-y|^{n}}  |f(y)|  dy  
=
 \int_{\Rn} \frac{|1 - (|x|/|y|)^{\alpha}|}{|x-y|^{n}} |y|^{\alpha}  |f(y)|  dy  ,
\end{split}
\end{equation}
so that by using Lemma \ref{SteinMainLemma} with $\phi = |\cdot| ^{\alpha} f$ we obtain
\begin{equation}
\|  T(|x|^{\alpha} f)  -   |x|^{\alpha} T f  \|_{L^{p}_{|x|}L^{\p}_{\theta}}
\lesssim \| |x|^{\alpha} f\|_{L^{p}_{|x|}L^{\p}_{\theta}} .
\end{equation}
Then, the desired estimate   
\eqref{eq:OurSteinWeiss} follows by (\ref{SWQuasiI}, \ref{SWQuasiII}) and triangle inequality.
Thus it only remains to prove Lemma \ref{SteinMainLemma}.

The idea is to use a change of variables which resembles the standard polar coordinates. 
In this variant the integration over the sphere is replaced by integration over the special orthogonal group $SO(n)$ and the
radial integration is replaced by integration over the multiplicative group of the positive real numbers. 
This method works efficiently 
when homogeneous power weights are involved; see e.g. \cite{DL, DenapoliDrelichmanDuran11-a}.

\subsubsection*{Proof of Lemma \ref{SteinMainLemma}}

By using the isomorphism 
\begin{equation*}
  \mathbb{S}^{n-1}\simeq SO(n)/SO(n-1)
\end{equation*}
we can rewrite integrals on $\mathbb{S}^{n-1}$ as follows
\begin{equation*}
  \int_{\mathbb{S}^{n-1}}g(y)dS(y)
  \simeq
  \int_{SO(n)}g(Ae)dA,\qquad n\ge2,
\end{equation*}
where $dA$ is the left Haar measure on
$SO(n)$, and $e\in\mathbb{S}^{n-1}$ is a fixed 
unit vector. Thus, via polar coordinates, 
a generic integral can be rewritten as
\begin{equation*}
\begin{split}
  \int_{\Rn} F(x,y) \phi(x) dy
   &=
  \int_{0}^{\infty}
  \int_{\mathbb{S}^{n-1}}
     F(x, \rho \omega)\phi(\rho \omega)dS_{\omega}\rho^{n-1}d \rho
    \\
  &\simeq
  \int_{0}^{\infty}
  \int_{SO(n)}F(x, \rho Be)\phi(\rho Be)dB\rho^{n-1}d \rho.
\end{split}
\end{equation*}
Hence the $L^{\p}_{\theta}$ norm can be written as
\begin{equation*}
\begin{split}
  & \bigg\| \int_{\Rn} F(|x|\theta,y) \phi(y) dy\bigg\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})}
  \simeq
  \bigg\| \int_{\Rn} F(|x| Ae,y) \phi(y) dy \bigg\|_{L^{\p}_{A}(SO(n))}
    \\
  &\le 
  \int_{0}^{\infty}
  \bigg\|
    \int_{SO(n)}F(|x|Ae, \rho Be)\phi(\rho Be)dB
  \bigg\|_{L^{\p}_{A}(SO(n))}
  \rho^{n-1}d \rho
\end{split} 
\end{equation*}
where $e$ is any fixed unit vector.
We choose $F$ as in \eqref{eq:SteinWeissKernel}
and we change variables
$B\to AB^{-1}$ in the inner integral.
By the invariance of the measure 
this is equivalent to
\begin{equation*}
\begin{split}
  & =\int_{0}^{\infty}
  \bigg\|
    \int_{SO(n)}\frac{|1 - (|x|/\rho)^{\beta}|}{|AB^{-1}(|x|Be-\rho e)|^n}\phi(\rho AB^{-1}e)dB
  \bigg\|_{L^{\p}_{A}(SO(n))}
  \rho^{n-1}d \rho
\\ 
&
=\int_{0}^{\infty}
  \bigg\|
    \int_{SO(n)}\frac{|1 - (|x|/\rho)^{\beta}|}{||x|Be-\rho e|^n}\phi(\rho AB^{-1}e)dB
  \bigg\|_{L^{\p}_{A}(SO(n))}
  \rho^{n-1}d \rho.
\end{split}
\end{equation*}
Notice that the integral
\begin{equation*}
    \int_{SO(n)}\frac{|1 - (|x|/\rho)^{\beta}|}{||x|Be-\rho e|^n}  \  |\phi(\rho AB^{-1}e)|dB
  =
  G*\phi(A)
\end{equation*}
is a convolution on $SO(n)$ of the functions
\begin{equation*}
  G(A)=\frac{|1 - (|x|/\rho)^{\beta}|}{||x|Ae-\rho e|^n},
  \qquad
  H(A)=|\phi(\rho Ae)|.
\end{equation*}
We can thus apply Young's inequality on $SO(n)$
(see for instance  
\cite[Theorem 1.2.12]{Grafakos08-a}) to obtain, for any
$1 \leq \p \leq \infty$,
the estimate
\begin{equation}\label{eq:firstest}
\begin{split}
 & \left\|\int_{\Rn}F(|x|\theta,y)\phi(y) dy\right\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})}
 \!\!\!
 \lesssim
  \int_{0}^{\infty}
  \left\| \frac{|1 - (|x|/\rho)^{\beta}|}{|(|x|e-\rho \theta)|^n} \right\|_{L^{1}_{\theta}(\mathbb{S}^{n-1})}
  \!\!\!\!\!\!\!\!\!\!\!\!\!
  \|\phi(\rho\theta)\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})}
  \rho^{n-1}d \rho
  \\ 
 & = 
  \int_{0}^{\infty}
  \bigg\| \frac{|1 - (|x|/\rho)^{\beta}|}{|(\frac{|x|}{\rho} -  \theta)|^n} \bigg\|_{L^{1}_{\theta}(\mathbb{S}^{n-1})}
  \!\!\!\!\!\!\!\!\!\!\!\!    \|\phi(\rho\theta)\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})}
 \frac{d \rho}{\rho}
  \end{split}
\end{equation}
where we switched back to the coordinates of $\mathbb{S}^{n-1}$.
Then we notice
\begin{equation}\label{est}
\begin{split}
\left \| \int_{\Rn} F(x,y) \phi(y)  dy  \right\|_{L^{p}_{|x|}L^{\p}_{\theta}} 
& 
\lesssim
\\ 
 &
\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!
 \lesssim 
 \Bigg\| \bigg\| \int_{\Rn} F(|x| \theta,y) \phi(y) dy  \bigg\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})} \!\!\!\!\!\!\!\!\!\!    dS_{\theta} |x|^{\frac np}  \Bigg\|_{L^{p}\left(\Rpiu(\cdot), d|x|/|x|\right)}
\end{split}
\end{equation}
where $\Rpiu(\cdot)$ is the moltiplicative group of positive real numbers equipped with its Haar measure $d\rho/\rho $.
Using \eqref{eq:firstest} allows to estimate \eqref{est} with
\begin{equation}\nonumber
\lesssim 
  \Bigg\|  
  \int_{0}^{\infty}
\left( \frac{|x|}{\rho} \right)^{\frac np}  
\bigg\|
\frac{|1 - (|x|/\rho)^{\beta}|}{\big| \big(\frac{|x|}{\rho} e - \theta \big) \big|^n} \bigg\|_{L^{1}_{\theta}(\mathbb{S}^{n-1})}
\!\!\!\!\!\!\!\!\!\!\!
\rho^{\frac np}
\|\phi(\rho\theta)\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})}
 \frac{d \rho}{\rho}
\Bigg\|_{L^{p}\left(\Rpiu(\cdot), d|x|/|x|\right)}.
\end{equation}
Notice that the inner term is a convolution on 
$\Rpiu (\cdot)$ of the functions
\begin{equation}\nonumber
g(\rho) = \rho^{\frac np}  
\left\| \frac{|1 - \rho^{\beta}|}{| (\rho e - \theta )|^n} \right\|_{L^{1}_{\theta}(\mathbb{S}^{n-1})},
\qquad
h(\rho) =\rho^{\frac np}
\|\phi(\rho\theta)\|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})}.
\end{equation}
Thus we can apply again Young's inequality to estimate further (\ref{est}) with
\begin{equation}\label{eq:SecondYoung}
 \lesssim
\big\|  \|  g(\rho) \|_{L^{1}_{\theta}(\mathbb{S}^{n-1})} \big\|_{L^{1}(\Rpiu (\cdot), d \rho / \rho)} 
\big\|  \|  h(\rho) \|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})} \big\|_{L^{p}(\Rpiu (\cdot), d \rho / \rho)},
\end{equation}
for all $1 \leq p \leq \infty$.
Once we have noticed that
\begin{equation}\nonumber
\big\|  \|  h(\rho) \|_{L^{\p}_{\theta}(\mathbb{S}^{n-1})} \big\|_{L^{p}(\Rpiu (\cdot), d \rho / \rho)}
= \| \phi \|_{L^{p}_{|x|}L^{\p}_{\theta}},
\end{equation}
the concluding step of the proof of the Lemma is represented by showing that the first term of \eqref{eq:SecondYoung} is bounded.
We split the integral 
\begin{equation}\nonumber
\int_{0}^{+\infty} \rho^{\frac np}  \int_{\mathbb{S}^{n-1}} \frac{|1 - \rho^{\beta}|}{|\rho e -\theta|^{n}} \  dS_{\theta}  \  \frac{d \rho}{\rho}
= \int_{0}^{\frac 12} (\cdot)
+
\int_{\frac 12}^{2} (\cdot)
+
\int_{2}^{+\infty} (\cdot)
=: I + II + III
\end{equation} 
and we bound separately the three terms.

If $0< \rho < 1/2$, then $| \rho e - \theta | \geq 1/2 $. Thus, since $|1 - \rho^{\beta}| < 1 + \rho^{\beta}$ 
\begin{equation}\nonumber
I \lesssim \int_{0}^{\frac 12} ( \rho^{\frac np -1} + \rho^{\frac np - 1 + \beta} ) d \rho < \infty
\end{equation}
provided that $p < \infty$ and $\beta > - n/p$.

If $1/2 \leq \rho \leq 2$, we notice that $|1 - \rho^{\beta}| \lesssim | 1-\rho |$ and we use that
(see for instance Lemma 2.1 in \cite{DL})
\begin{equation}\nonumber
\int_{\mathbb{S}^{n-1}} |\rho e - \theta|^{-n} dS_{\theta}
 \simeq \frac{1}{|1- \rho |}
\end{equation} 
to bound
\begin{equation}
II \simeq
\int_{\frac 12}^{2} \rho^{\frac np -1} \frac{|1 - \rho^{\beta}|}{|1- \rho|} d \rho 
\lesssim
\int_{\frac 12}^{2} \rho^{\frac np -1}  d \rho
<
\infty.
\end{equation}

Finally, if $2 < \rho < + \infty$, then $| \rho e - \theta | \geq | \rho | - | \theta | \geq | \rho | /2 $. Thus, since $|1 - \rho^{\beta}| < 1 + \rho^{\beta}$ 
\begin{equation}\nonumber
III \lesssim \int_{2}^{+ \infty} ( \rho^{\frac np -1 -n} + \rho^{\frac np - 1 + \beta - n} ) d \rho < \infty
\end{equation}
provided that $p > 1$ and $\beta < n - n/p$,
that concludes the proof. 

\hfill $\Box$

{\bf Acknowledgements.} The first author is supported by the FIRB 2012 \lq Dispersive dynamics, Fourier analysis and variational methods'. Part of the work was done during the first author's visit at MSRI, Berkeley (CA), within the 
program \lq New Challenges in PDE: Deterministic Dynamics and Randomness in High and Infinite Dimensional Systems', which he acknowledges for the wonderful working conditions. 

The second author is supported by the
ERC grant 277778 and MINECO grant SEV-2011-0087 (Spain).

\end{document}